Determine the lines tangent to the circle x^2+y^2-2x-8y+15=0 and x-y+10=0. Sketch the graph
Found 2 solutions by Alan3354, ikleyn:
Answer by Alan3354(69443)

Determine the lines tangent to the circle x^2+y^2-2x-8y+15=0 and x-y+10=0. Sketch the graph
=======================
x^2+y^2-2x-8y+15=0 is a circle.
-----------
IDK what you mean by tangent to x-y+10=0.

Answer by ikleyn(52879)

.

The problem is  POORLY  stated.

So poorly,  that the post  CAN  NOT  be considered as a  Math problem.